Solve by substitution method.
6x-7y=-35
2x+21=y
Solution type in ordered pair.

Answer by mananth(16949)

6x-7y=-35 .............1
y=21+2x ---------2
Plug the value of y in (1)
6x-7(21+2x)=-35
6x-147-14x=-35
6x-14x=-35+147
-8x=112
/-8
x=-14
Plug the value of x in (1)
6x-7y=-35
6*-14-7y=-35
-84-7y= -35
-7y=49
/-7
y=-7